OpenStudy (anonymous):

what is the defining equation of the reciprocal of y= 3/(2x-1)

OpenStudy (anonymous):

switch x=y and y=x y=3/(2x-1) x=3/(2y-1) solve for y=____

OpenStudy (anonymous):

@BabiiMuffin wouldnt that be the inverse..?

OpenStudy (anonymous):

oh woops reciprocal is just switch fractions upsidedown

OpenStudy (anonymous):

oh right. hahah yeah i got it thanks @BabiiMuffin ! :)